# Principal component analysis balancing prediction and approximation accuracy for spatial data

### v1

@@ 1 Introduction
Principal component analysis (PCA) is a classical dimension-reduction technique leading to independent and lower-dimensional scores, called principal component scores or PC scores, approximating the multivariate measurements (Jolliffe, 2002). Existing methodologies that tackle some of the aforementioned challenges are often based on extensions of PCA. Jandarov et al. (2017) proposed a spatially predictive PCA algorithm, where the PC scores were constrained to be linear combinations of some covariates and spatial basis, so they can be more accurately predicted at new locations. Bose et al. (2018) extended this predictive PCA approach and enabled adaptive selection of covariates to be included for each PC. Vu et al. (2020, 2022) proposed a probabilistic version of predictive PCA along with a low-rank matrix completion algorithm, which offers improved performance when there is spatially informative missing data. Keller et al. (2017) developed a predictive $k$-means algorithm where dimension reduction was conducted by clustering.
@@
In spatial transcriptomics applications such as domain detection (see detailed discussion in Section 5.2), Zhao et al. (2021) processed gene expression information using PCA and conducted downstream clustering via a Bayesian approach, where the spatial arrangement of the measured spots were modeled using the PC scores. Shang and Zhou (2022) proposed a probabilistic PCA algorithm, where spatial information was incorporated into a latent factor model for gene expression. Liu et al. (2022) developed a joint approach that simultaneously conducted dimension reduction and clustering, and used a latent hidden Markov random field model to enforce spatial structures of the clusters.
@@
When dimension reduction is used as an initial step before downstream analyses such as prediction, statistical inference and clustering, there are two key considerations for the quality of dimension reduction: The first is representability, which is how well the lower dimensional components approximate the original measurements. The second, often conflicting, goal is predictability of the resulting components, so that they preserve meaningful scientific and spatial information when extrapolated to locations without measurements; such predictability ensures that the extrapolations are of high quality for subsequent modeling. Even in analytical tasks that do not primarily focus on prediction (e.g., domain detection in spatial transcriptomics), the predictability of the PC scores is still desirable and implicitly considered, since it enforces the interpretability and spatial structure of the PCs; see Section 5.2 for further illustration.
@@
Among existing methodologies, classical PCA solely focuses on representability, while predictive PCA (Jandarov et al., 2017) and its variants (Bose et al., 2018) prioritize predictability and optimize representability only after the former is guaranteed. Probabilistic approaches such as Shang and Zhou (2022), Liu et al. (2022) and Vu et al. (2020) implicitly incorporate both tasks into the latent factor models, though their performance depends on the validity of the parametric model assumptions, and there is no explicit interpretation or optimality guarantee on either property or the trade-off between them.
@@
In this work, we propose a flexible dimension reduction algorithm, termed representative and predictive PCA (RapPCA), that explicitly minimizes a combination of representation and prediction errors and finds the optimal balance between them. We further allow the underlying lower-dimensional scores to have complex spatial structure and non-linear relationships with external covariates (if any). We show that the optimization problem involved can be solved by eigendecomposition on transformed data, enabling simple and efficient computation.
@@ 2.1 Classical and Predictive PCA
Our ultimate goal is to conduct prediction, clustering or other modeling tasks on $Y$ based on the spatial coordinates and covariates. Since the columns of $Y$ may be strongly correlated and/or noisy (Rao et al., 2021), dimension reduction can be used to to extract scientifically meaningful signals from the original data. The classical PCA achieves this by maximizing the proportion of variability in $Y$ that is explained by the lower dimensional PCs. We briefly introduce the classical PCA algorithm under the formulation of Shen and Huang (2008) to highlight its connection with related techniques.
@@
$\min_{u_{l},v_{l}}\ \left\lVert Y^{(l)}-u_{l}v_{l}^{\top}\right\rVert_{F}^{2}\quad\text{s.t. }\lVert v_{l}\rVert_{2}=1$ (1)
where the 1-dimensional PC score $u_{l}$ is an $n$-vector, the loading $v_{l}$ is $p\times 1$, and $\lVert\cdot\rVert_{F}$ and $\lVert\cdot\rVert_{2}$ represent Frobenius norm for matrices and $\ell_{2}$ norm for vectors, respectively. $Y^{(l)}$ is the residual from approximation after each iteration. In other words, denoting $\tilde{u}_{l},\tilde{v}_{l}$ as the solution to (1), we define $Y^{(l)}=Y^{(l-1)}-\tilde{u}_{l-1}\tilde{v}_{l-1}^{\top}$ for $2\leq l\leq r$ and $Y^{(1)}=Y$.
Combining the optimal solutions as $\tilde{U}=[\tilde{u}_{1},\ldots,\tilde{u}_{r}]$, we obtain the $r$-dimensional PC scores $\tilde{U}$ as an approximation of the $p$-dimensional data $Y$. However, while each of the $u_{l}$’s explain the greatest variability in $Y$, there is no guarantee that they are scientifically interpretable, and they do not explicitly account for the spatial structures underlying the observations in $Y$. A natural idea is to constrain each PC score $u_{l}$ within certain model space based on the spatial coordinates and covariates $X$ (if any), which motivates the predictive PCA algorithm (Jandarov et al., 2017).
Predictive PCA builds upon the expression in (1) and solves
@@
for each $l=1,\ldots,r$. Here, $Z=[X\ B]$, where $B\in\mathbb{R}^{n\times m}$ contains $m$ thin-plate regression spline basis functions capturing the spatial effects, and in contrast to (1), normalization is done on the term $Z{\alpha_{l}}$ instead of $v_{l}$. Compared with (1), this formulation restricts each PC score $u_{l}=\frac{Z\alpha_{l}}{\lVert Z\alpha_{l}\rVert_{2}}$ to fall within a certain model space, and more specifically, the space spanned by the columns of $Z$. $Y^{(l)}$ is defined as the residual after approximation, i.e. $Y^{(l)}=Y^{(l-1)}-Y^{(l-1)}\frac{\tilde{v}_{l-1}\tilde{v}_{l-1}^{\top}}{\lVert\tilde{v}_{l-1}\rVert_{2}}$ for $2\leq l\leq r$ and $Y^{(1)}=Y$. Here, the quantity $Y^{(l-1)}\tilde{v}_{l-1}$ corresponds to $\tilde{u}_{l-1}$ as in the expression of residuals for PCA.
After solving either (1) or (2) for all $r$ PCs, the loadings are $\ell_{2}$-normalized ($\tilde{v}_{r}\leftarrow\tilde{v}_{r}/\lVert\tilde{v}_{r}\rVert_{2}$) if not already done, and they are concatenated as $\tilde{V}_{p\times r}:=[\tilde{v}_{1}\ \ldots\ \tilde{v}_{r}]$. The PC scores are then defined as $\tilde{U}_{n\times r}:=Y\tilde{V}$.
By restricting the induced PC score to fall within the linear span of $Z$, the solution to (2) forces the resulting PC score to have spatial smoothness and contain information from the covariates $X$. This solution is therefore more predictable by $X$ along with spatial smoothing, compared to classical PCA.
As noted earlier, classical PCA aims to achieve optimal representability of the PC scores $\tilde{u}$, but these $\tilde{u}$’s may not retain meaningful signals in $X$ or important spatial patterns to be well predictable. In contrast, predictive PCA and its variants (Jandarov et al., 2017; Bose et al., 2018) minimize the approximation gap, which will be formally defined in Section 2.2, after constraining each $\tilde{u}$ to lie within a specific model space; but when these PC scores $\tilde{u}$ are predicted at new locations and transformed back to the higher dimensional space of $Y$, closeness in the space of $\tilde{u}$ may not necessarily translate to the original space of observations $Y$ due to reduced approximation accuracy. Our proposed method is based on an interpolation between classical and predictive PCA, and encourages $\tilde{u}$ to be close to, but not exactly within, some model space while balancing the quality of representation. This proposal will be introduced in detail in Section 3.
@@ 2.2 Evaluation Metrics for Dimension Reduction
Before introducing our proposal, it is helpful to formalize different aspects of dimension reduction performances, such as representability and predictability, and to characterize the overall balance between them. Though the ultimate goal may not exactly be the prediction of $Y$, we quantify the predictability of the PC scores in the training-test setting because desirable predictability indicates important scientific information (based on the external covariates $X$) and spatial structure being preserved in the resulting PC scores.
Suppose the PC scores $\tilde{U}_{\text{trn}}=[\tilde{u}_{1}\ \ldots\tilde{u}_{r}]$ and loadings $\tilde{V}$ are obtained on a set of training data, and $\tilde{u}_{1},\ldots,\tilde{u}_{r}$ are then predicted at the test locations with a user-specified spatial prediction model as $\hat{U}_{\text{tst}}:=[\hat{u}_{1}\ \ldots\ \hat{u}_{r}]$. We define the mean-squared prediction error (MSPE) of this procedure as
$\text{MSPE}=n^{-1}\left\lVert(\hat{U}_{\text{tst}}-U^{*})\tilde{V}^{\top}\right\rVert_{F}^{2}$
@@
$\text{MSRE}=n^{-1}\left\lVert Y_{\text{tst}}-U^{*}\tilde{V}^{\top}\right\rVert_{F}^{2},$
@@
Given any outcome data $Y$, loadings $\tilde{V}$ from some dimension reduction procedure and predicted PC scores $\hat{U}$, we define the total mean-squared error (TMSE) resulting from both dimension reduction and prediction as $\text{TMSE}=n^{-1}\left\lVert Y-\hat{U}\tilde{V}^{\top}\right\rVert_{F}^{2}$. It measures the discrepancy between the true data $Y$ and the predicted scores that are transformed back to the original, higher-dimensional space. In the specific training-test setting, the expression of TMSE becomes
$\text{TMSE}=n^{-1}\left\lVert Y_{\text{tst}}-\hat{U}_{\text{tst}}\tilde{V}^{\top}\right\rVert_{F}^{2},$
and hence MSPE and MSRE can be viewed informally as a decomposition of the overall TMSE, where the prediction and approximation gaps, i.e., $(U^{*}-\hat{U}_{\text{tst}})\tilde{V}^{\top}$ and $Y_{\text{tst}}-U^{*}\tilde{V}^{\top}$, are components of the overall error $Y_{\text{tst}}-\hat{U}_{\text{tst}}\tilde{V}^{\top}$.
@@ 3 Representative and Predictive PCA (RapPCA)
Noting that the model space can be generalized from the linear span in (2) to incorporate more flexible mechanisms underlying each PC $\tilde{u}$, we adopt kernel methods (Evgeniou et al., 2000; Hastie et al., 2009) to capture non-linear covariate effects; specifically, we use smoothing splines (Wahba, 1990; Wood, 2003) rather than unpenalized regression splines to enforce spatial structures.
Combining these ideas, we propose to solve the following sequence of optimization problems, for $l=1,\ldots,r$, to extract $r$ PCs from $Y$:
@@
Here, $Y^{(l)}$ is defined similarly to the notation in Section 2 as the residual $Y^{(l-1)}-\tilde{u}_{l-1}{\tilde{v}}_{l-1}^{\top}$ for $2\leq l\leq r$, and set to be the original data $Y$ when $l=1$. Moreover, $K$ is a kernel matrix such that $K_{ij}=k(X_{i\cdot},X_{j\cdot})$ for some kernel function $k(\cdot,\cdot)$, with $X_{i\cdot}$ being the $i$th observation (row) of $X$. The columns of $B$ are evaluations of $m$ spline basis functions, e.g., thin-plate regression splines as in predictive PCA, at the spatial coordinates $\{(s_{i1},s_{i2})\}_{i=1}^{n}$, and $Q$ is the penalty matrix induced by the choice of smoothing splines. For example, the penalty $Q$ corresponding to the thin-plate regression splines is the wiggliness penalty defined in Wood (2003); and for P-splines, the penalty term depends on the differences between consecutive $\beta_{k}$’s, i.e., $\sum_{k}(\beta_{k}-\beta_{k-1})^{2}$ (Wood, 2017). The computation of this optimization problem could suffer when $K$ or $Q$ is near-singular. Therefore in practice, we replace $K$ and $Q$ with $\tilde{K}:=K+\delta I_{n}$ and $\tilde{Q}:=Q+\delta I_{m}$ for a small, data-independent constant $\delta$ and identity matrices $I_{n},I_{m}$ in the penalty terms to avoid near-singularity; see also Li et al. (2019) for similar handling. This leads to
@@
In (3), the first term $\left\lVert Y^{(l)}-uv^{\top}\right\rVert_{F}^{2}$ measures the approximation gap of the score $u$ and loading $v$, similar to the objective of minimization for classical PCA. The second term reflects the distance between the score $u$ and the model space specified by $K$ and $B$, and encourages $u$ to be predictable based on the covariates and spatial coordinates. In particular, the kernel approach allows for non-linear relationships between $X$ and $u$, and is more flexible than most existing, predictive variants of PCA (e.g., Jandarov et al., 2017; Bose et al., 2018; Vu et al., 2020; Shang and Zhou, 2022) which specify linear relationships. The constraint $u=Y^{(l)}v$ ensures a similar relationship between the PC scores and loadings as in classical PCA. The tuning parameter $\gamma$ controls the trade-off between predictability and representability, with larger $\gamma$ imposing higher penalty on unpredictable scores $u$. The third and fourth terms in (3) are $\ell_{2}$ penalty terms enforcing the identifiability of the model parameters and helping to avoid overfitting the data.
Though optimizing $v,\alpha,\beta$ iteratively via coordinate descent is straightforward in implementation, the optimization problem (3) is non-convex due to the joint optimization of $v$ and $(\alpha,\beta)$, as well as the non-convex feasible set of $v$. Consequently, coordinate gradient descent is not guaranteed to converge to the global minimizer. Instead, we propose an analytical solution to (3) that attains the global miminum despite the non-convexity of the objective function $f_{\gamma,\lambda_{1},\lambda_{2}}(v,\alpha,\beta)$ and/or the constraints. We describe our proposed procedure in Algorithm 1, and refer to the proposed method as representative and predictivePCA, or RapPCA.
@@
Input: Number of PCs to extract $r>0$; hyper-parameters $\gamma^{(1)},\ldots,\gamma^{(l)}$, $\lambda_{1}^{(1)},\ldots,\lambda_{1}^{(l)}$, $\lambda_{2}^{(1)},\ldots,\lambda_{2}^{(l)}$ specified for each PC; kernel function $k(\cdot,\cdot)$; spatial spline basis evaluated at $\{(s_{i1},s_{i2})\}$: $B^{n\times m}$; penalty matrix $Q_{m\times m}$; $\delta>0$

### v2

@@ 1 Introduction
Principal component analysis (PCA) is a classical dimension-reduction technique. It leads to independent and lower-dimensional scores, called principal component scores, or PC scores, that approximate the multivariate measurements (Jolliffe, 2002). Existing methodologies that tackle some of the aforementioned challenges are often based on extensions of PCA. Jandarov et al. (2017) proposed a spatially predictive PCA algorithm, where the PC scores were constrained to be linear combinations of covariates and spatial basis, so they can be more accurately predicted at new locations. Bose et al. (2018) extended this predictive PCA approach and enabled adaptive selection of covariates to be included for each PC. Vu et al. (2020, 2022) proposed a probabilistic version of predictive PCA along with a low-rank matrix completion algorithm, which offers improved performance when there is spatially informative missing data. Keller et al. (2017) developed a predictive $k$-means algorithm where dimension reduction was conducted by clustering.
@@
In spatial transcriptomics applications such as domain detection (see detailed discussion in Section 5.2), Zhao et al. (2021) processed gene expression information using PCA and conducted downstream clustering via a Bayesian approach, where the spatial arrangement of the measured spots were modeled using the PC scores. Shang and Zhou (2022) proposed a probabilistic PCA algorithm, where spatial information was incorporated into a latent factor model for gene expression. Liu et al. (2022) developed a joint approach that simultaneously performs dimension reduction and clustering, and uses a latent hidden Markov random field model to enforce spatial structures of the clusters.
@@
When dimension reduction is used as an initial step before downstream analyses such as prediction, statistical inference and clustering, there are two key considerations for the quality of dimension reduction. The first is representability, which is how well the lower dimensional components approximate the original measurements. The second, often conflicting, goal is predictability of the resulting components, so that they preserve meaningful scientific and spatial information when extrapolated to locations without measurements; such predictability ensures that the extrapolations are of high quality for subsequent modeling. Even in analytical tasks that do not primarily focus on prediction (e.g., domain detection in spatial transcriptomics), the predictability of the PC scores is still desirable and implicitly considered, since it enforces the interpretability and spatial structure of the PCs; see Section 5.2 for further illustration.
@@
Among existing methodologies, classical PCA solely focuses on representability, while predictive PCA (Jandarov et al., 2017) and its variants (Bose et al., 2018) prioritize predictability and optimize representability only after the former is guaranteed. Probabilistic approaches, such as Shang and Zhou (2022), Liu et al. (2022) and Vu et al. (2020), implicitly incorporate both tasks into the latent factor models, though their performance depends on the validity of the parametric model assumptions, and there is no explicit interpretation or optimality guarantee on either property or the trade-off between them.
@@
In this work, we propose a flexible dimension reduction algorithm, termed representative and predictive PCA (RapPCA), that explicitly minimizes a combination of representation and prediction errors and finds the optimal balance between them. We further allow the underlying lower-dimensional scores to have complex spatial structure and non-linear relationships with external covariates (if any). We show that the optimization problem involved can be solved by eigendecomposition of transformed data, enabling simple and efficient computation.
@@ 2.1 Classical and Predictive PCA
The ultimate goal is to conduct prediction, clustering or other modeling tasks on $Y$ based on the spatial coordinates and covariates. Since the columns of $Y$ may be strongly correlated and/or noisy (Rao et al., 2021), dimension reduction can be used to to extract scientifically meaningful signals from the original data. The classical PCA achieves this by maximizing the proportion of variability in $Y$ that is explained by the lower dimensional PCs. We briefly introduce the classical PCA algorithm under the formulation of Shen and Huang (2008) to highlight its connection with related techniques.
@@
$\min_{u_{l},v_{l}}\ \left\lVert Y^{(l)}-u_{l}v_{l}^{\top}\right\rVert_{F}^{2}\quad\text{s.t. }\lVert v_{l}\rVert_{2}=1,$ (1)
where the 1-dimensional PC score $u_{l}$ is an $n$-vector, the loading $v_{l}$ is $p\times 1$, and $\lVert\cdot\rVert_{F}$ and $\lVert\cdot\rVert_{2}$ represent Frobenius norm for matrices and $\ell_{2}$ norm for vectors, respectively. $Y^{(l)}$ is the residual from approximation after each iteration: denoting $\tilde{u}_{l},\tilde{v}_{l}$ as the solution to (1), we define $Y^{(l)}=Y^{(l-1)}-\tilde{u}_{l-1}\tilde{v}_{l-1}^{\top}$ for $2\leq l\leq r$ and $Y^{(1)}=Y$.
While each of the $u_{l}$’s explain the greatest variability in $Y$, there is no guarantee that they are scientifically interpretable, and they do not explicitly account for the spatial structures underlying $Y$. Recognizing this, predictive PCA (Jandarov et al., 2017) builds upon the expression in (1), but constrains each PC score $u_{l}$ to be in a model space of the spatial coordinates and covariates $X$, leading to the optimiztion problem,
@@
for each $l=1,\ldots,r$. Here, $Z=[X\ B]$, $B\in\mathbb{R}^{n\times m}$ contains $m$ thin-plate regression spline basis functions capturing the spatial effects, and in contrast to (1), normalization is done on the term $Z{\alpha_{l}}$ instead of $v_{l}$. As before, $Y^{(l)}$ is the residual after approximation, i.e. $Y^{(l)}=Y^{(l-1)}-Y^{(l-1)}\frac{\tilde{v}_{l-1}\tilde{v}_{l-1}^{\top}}{\lVert\tilde{v}_{l-1}\rVert_{2}}$ for $2\leq l\leq r$ and $Y^{(1)}=Y$. Note that the quantity $Y^{(l-1)}\tilde{v}_{l-1}$ corresponds to $\tilde{u}_{l-1}$ in the expression of PCA. Compared with (1), this formulation restricts each PC score $u_{l}=\frac{Z\alpha_{l}}{\lVert Z\alpha_{l}\rVert_{2}}$ to fall within the space spanned by the columns of $Z$. This ensures the spatial smoothness and the inclusion of information from covariates $X$ in the PC score $u_{l}$, and thus enforces predictability.
After solving either (1) or (2) for all $r$ PCs, the loadings are $\ell_{2}$-normalized ($\tilde{v}_{r}\leftarrow\tilde{v}_{r}/\lVert\tilde{v}_{r}\rVert_{2}$) if not already done, and then concatenated as $\tilde{V}_{p\times r}:=[\tilde{v}_{1}\ \ldots\ \tilde{v}_{r}]$. The PC scores are then defined as $\tilde{U}_{n\times r}:=Y\tilde{V}$.
As noted earlier, classical PCA aims to achieve optimal representability of the PC scores $\tilde{u}$, but these $\tilde{u}$’s may not retain meaningful signals in $X$ or important spatial patterns to be well predictable. In contrast, predictive PCA and its variants (Jandarov et al., 2017; Bose et al., 2018) minimize the approximation gap, which will be formally defined in Section 2.2, after constraining each $\tilde{u}$ to lie within a specific model space; but when these PC scores $\tilde{u}$ are predicted at new locations and transformed back to the higher dimensional space of $Y$, closeness in the space of $\tilde{u}$ may not necessarily translate to the original space of observations $Y$ due to reduced approximation accuracy. Our proposed method, introduced in Section 3, is based on an interpolation between classical and predictive PCA, and encourages $\tilde{u}$ to be close to, but not exactly within, the specified model space while balancing the quality of representation.
@@ 2.2 Evaluation Metrics for Dimension Reduction
Before introducing our proposal, it is helpful to formalize different aspects of dimension reduction performances, such as representability and predictability, and to characterize the overall balance between them.
Suppose the PC scores $\tilde{U}_{\text{trn}}=[\tilde{u}_{1}\ \ldots\tilde{u}_{r}]$ and loadings $\tilde{V}$ are obtained on a set of training data with size $n_{\text{trn}}$, and $\tilde{u}_{1},\ldots,\tilde{u}_{r}$ are then predicted at $n_{\text{tst}}$ test locations with a user-specified spatial prediction model as $\hat{U}_{\text{tst}}:=[\hat{u}_{1}\ \ldots\ \hat{u}_{r}]$. We define the mean-squared prediction error (MSPE) of this procedure as
$\text{MSPE}={n_{\text{tst}}}^{-1}\left\lVert(\hat{U}_{\text{tst}}-U^{*})\tilde{V}^{\top}\right\rVert_{F}^{2},$
@@
$\text{MSRE}={n_{\text{tst}}}^{-1}\left\lVert Y_{\text{tst}}-U^{*}\tilde{V}^{\top}\right\rVert_{F}^{2},$
@@
Given any outcome data $Y$, loadings $\tilde{V}$ from a dimension reduction procedure and predicted PC scores $\hat{U}$, we define the total mean-squared error (TMSE) resulting from both dimension reduction and prediction as $\text{TMSE}=n^{-1}\left\lVert Y-\hat{U}\tilde{V}^{\top}\right\rVert_{F}^{2}$. This quantity measures the discrepancy between the true data $Y$ and the predicted scores that are transformed back to the original, higher-dimensional space. In the specific training-test setting, the expression of TMSE becomes
$\text{TMSE}={n_{\text{tst}}}^{-1}\left\lVert Y_{\text{tst}}-\hat{U}_{\text{tst}}\tilde{V}^{\top}\right\rVert_{F}^{2}.$
MSPE and MSRE can be viewed informally as a decomposition of the overall TMSE, where the prediction and approximation gaps, i.e., $(U^{*}-\hat{U}_{\text{tst}})\tilde{V}^{\top}$ and $Y_{\text{tst}}-U^{*}\tilde{V}^{\top}$, are components of the overall error $Y_{\text{tst}}-\hat{U}_{\text{tst}}\tilde{V}^{\top}$.
@@ 3 Representative and Predictive PCA (RapPCA)
Noting that the model space can be generalized from the linear span in (2) to incorporate more flexible mechanisms underlying each PC $\tilde{u}$, we adopt kernel methods (Evgeniou et al., 2000; Hastie et al., 2009) to capture non-linear covariate effects; additionally, we use smoothing splines (Wahba, 1990; Wood, 2003) rather than unpenalized regression splines to enforce spatial structures.
More conceretely, we propose to solve the following sequence of optimization problems, for $l=1,\ldots,r$, to extract $r$ PCs from $Y$:
@@
As in Section 2, $Y^{(l)}$ denotes the residual $Y^{(l-1)}-\tilde{u}_{l-1}{\tilde{v}}_{l-1}^{\top}$ for $2\leq l\leq r$, and is set to the original data $Y$ when $l=1$. Moreover, $K{\in\mathbb{R}^{n\times n}}$ is a kernel matrix such that $K_{ij}=k(X_{i\cdot},X_{j\cdot})$ for some kernel function $k(\cdot,\cdot)$, with $X_{i\cdot}$ being the $i$th observation (row) of $X$. The kernel function $k(\cdot,\cdot)$ maps the covariates $X$ into a higher-dimensional space, and the kernel matrix $K$ encodes the pairwise similarities between each entry $X_{i\cdot}$ in this transformed space. Linear operations in the high-dimensional space effectively captures non-linear relationships in the original input space of $X$. This enables more flexibility than most existing predictive variants of PCA which specify linear relationships (e.g., Jandarov et al., 2017; Bose et al., 2018; Vu et al., 2020; Shang and Zhou, 2022). Common choices of kernel functions include linear kernel, $k(x,x^{\prime})=x^{\top}x^{\prime}$, and Gaussian kernel, $k_{h}(x,x^{\prime})=\exp\left(-h\lVert x-x^{\prime}\rVert^{2}\right)$ (Hofmann et al., 2008; Murphy, 2012). The columns of $B{\in\mathbb{R}^{n\times m}}$ are evaluations of $m$ spline basis functions, e.g., thin-plate regression splines as in predictive PCA, at the spatial coordinates $\{(s_{i1},s_{i2})\}_{i=1}^{n}$, and $Q{\in\mathbb{R}^{m\times m}}$ is the penalty matrix induced by the choice of smoothing splines. For example, the penalty $Q$ corresponding to the thin-plate regression splines is the wiggliness penalty defined in Wood (2003); for P-splines, the penalty term depends on the differences between consecutive entries of $\beta$, i.e., $\sum_{k}(\beta_{k}-\beta_{k-1})^{2}$ (Wood, 2017).
In (3), the first term $\left\lVert Y^{(l)}-uv^{\top}\right\rVert_{F}^{2}$ measures the approximation gap of the score $u$ and loading $v$, similar to the objective of minimization for classical PCA. The second term reflects the distance between the score $u$ and the model space specified by $K$ and $B$, and encourages $u$ to be predictable based on the covariates and spatial coordinates. The constraint $u=Y^{(l)}v$ ensures a similar relationship between the PC scores and loadings as in classical PCA.
The computation of the optimization problem in (3) could suffer when $K$ or $Q$ is near-singular. Therefore in practice, we replace $K$ and $Q$ with $\tilde{K}:=K+\delta I_{n}$ and $\tilde{Q}:=Q+\delta I_{m}$ for a small, data-independent constant $\delta$, and identity matrices $I_{n}$ and $I_{m}$ in the penalty terms to avoid near-singularity; see also Li et al. (2019) for similar handling. This leads to
@@
Here, $\gamma$, $\lambda_{1}$ and $\lambda_{2}$ are tuning parameters, among which $\gamma$ is of most interest as it controls the trade-off between predictability and representability, with larger $\gamma$ imposing higher penalty on unpredictable scores $u:=Y^{(l)}v$. The third and fourth terms in (4) are $\ell_{2}$ penalties that enforce the identifiability of the model parameters and help to avoid overfitting the data.
Though optimizing $v,\alpha,\beta$ iteratively via coordinate descent is straightforward to implement, the optimization problem (4) is non-convex due to the joint optimization of $v$ and $(\alpha,\beta)$, as well as the non-convex feasible set of $v$. Consequently, iterative algorithms, including coordinate descent are not guaranteed to converge to the global minimizer. Instead, we propose an analytical solution to (4) that attains the global miminum despite the non-convexity of the objective function $f_{\gamma,\lambda_{1},\lambda_{2}}(v,\alpha,\beta)$ and/or the constraints. We describe our proposed procedure in Algorithm 1, and refer to the proposed method as representative and predictivePCA, or RapPCA. Theorem 1, proved in Appendix A, establishes the optimality of $\tilde{U}$ and $\tilde{V}$. Numerical examples in Appendix B.1 empirically verify the optimality.
@@
Input: Number of PCs to extract $r>0$; hyper-parameters $\gamma^{(l)}$, $\lambda_{1}^{(l)}$, $\lambda_{2}^{(l)}$ specified for the $l$th PC ($l=1,\ldots,r$); kernel function $k(\cdot,\cdot)$; spatial spline basis evaluated at $\{(s_{i1},s_{i2})\}$: $B^{n\times m}$; penalty matrix $Q_{m\times m}$; $\delta>0$